OpenStudy (anonymous):

how about 2x=3y+6 and 4x-3y+12?

OpenStudy (dumbcow):

is that 2x = 3y+6 4x = 3y+12 ??

OpenStudy (anonymous):

ok thats ryt thank u:)

OpenStudy (dumbcow):

then subtract the equations 2x = 3y +6 -4x = -3y -12 -------------- -2x = -6 x = 3 2(3) = 3y + 6 6=3y+6 3y = 0 y = 0 solution: (3,0)